OpenStudy (anonymous):

X+1/x-2 how to find the range of it?

OpenStudy (er.mohd.amir):

range means value of function when put the different value of x

OpenStudy (priyar):

what is 3/0? @Zaada ?

OpenStudy (anonymous):

No defined :)

OpenStudy (priyar):

good..its not defined..So the denominator can never be zero..right?

OpenStudy (priyar):

then what value should x not take..so that x-2 is not equal to 0

OpenStudy (anonymous):

2

OpenStudy (priyar):

great job!

OpenStudy (anonymous):

But, is not that how we find the domain of the function?

OpenStudy (priyar):

yes u r clear with that?

OpenStudy (anonymous):

Yeap!

OpenStudy (priyar):

to find the range(proper way) take this expression equal to some 'y' and rearrange the new eq to remove the denominator..ok?

OpenStudy (anonymous):

So, smth like this? : 2(x-2)=x-1 and solve for x?

OpenStudy (priyar):

how did u het this? where is 'y'?

OpenStudy (priyar):

*get

OpenStudy (priyar):

its more like.. \[y=\frac{ x+1 }{ x-2 }\] ok? now multiply with (x-2) on both sides

OpenStudy (priyar):

@Zaada ??

OpenStudy (anonymous):

oh, sorry. I was away for a min. So yes, I did multiply by (x-2) substituting y by 2.

OpenStudy (priyar):

hey..where is y? i don't see any 'y' on the left side..

OpenStudy (anonymous):

\[y(x-2)=x+1\] What do I do now?

OpenStudy (anonymous):

I see, so all real # except 1? @priyar

OpenStudy (priyar):

now multiply y inside..

OpenStudy (priyar):

and write in this form:x=..

OpenStudy (anonymous):

x=2y+1/y-1